OpenStudy (anonymous):

Essay; show all work. Bob’s Barber Shop estimates their gross revenue for the second quarter to be given by the polynomial 4x3 – 6x2 – 6x + 5. The shop estimates their costs for that quarter to be given by 4x2 – 7x + 5. For the second quarter, find and simplify a polynomial that will represent their profit. Note: Profit = Revenue - Costs

OpenStudy (agreene):

(4x^3 – 6x^2 – 6x + 5)-(4x^2 – 7x + 5) \[\large4x^3-10x^2+x\]

OpenStudy (anonymous):

tks thats what i came up with i just did it the long long way

OpenStudy (agreene):

I just didnt show all my work ;) there isnt a short way, that I know of.

OpenStudy (anonymous):

R(x) = 4*x^3 - 6*x^2 - 6*x + 5 C(x) = 4*x^2 - 7*x + 5 P(x) = R(x) - C(x) P(x) = 4*x^3 - 6*x^2 - 6*x + 5 - 4*x^2 + 7*x - 5 P(x) = 4*x^3 - 10*x^2 + x

OpenStudy (agreene):

Yeah, I just didnt write all that down :D

OpenStudy (anonymous):

tks i have one more then am gonna quit